# MSRS: Adaptive Multi-Subspace Representation Steering for Attribute Alignment in Large Language Models

### v3

@@ Abstract
Activation steering offers a promising approach to controlling the behavior of Large Language Models by directly manipulating their internal activations. However, most existing methods struggle to jointly steer multiple attributes, often resulting in interference and undesirable trade-offs. To address this challenge, we propose Multi-Subspace Representation Steering (MSRS), a novel framework for effective multi-attribute steering via subspace representation fine-tuning. MSRS reduces inter-attribute interference by allocating orthogonal subspaces to each attribute, isolating their influence within the model’s representation space. MSRS also incorporates a hybrid subspace composition strategy: it combines attribute-specific subspaces for unique steering directions with a shared subspace for common steering directions. A dynamic weighting function learns to efficiently integrate these components for precise control. During inference, MSRS introduces a token-level steering mechanism that dynamically identifies and intervenes on the most semantically relevant tokens, enabling fine-grained behavioral modulation. Experimental results show that MSRS significantly reduces attribute conflicts, surpasses existing methods across a range of attributes, and generalizes effectively to diverse downstream tasks.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, driving advancements in applications such as text generation, question answering, and dialogue systems (Qin et al., 2024; Matarazzo & Torlone, 2025; Su et al., 2023a; b; Cheng et al., 2024; Yang et al., 2024b; Zhang et al., 2024; Yao et al., 2025a). However, as LLMs are increasingly deployed in real-world, sensitive contexts, ensuring their behavior aligns with desired attributes, such as truthfulness and fairness, has become a critical challenge (Yang et al., 2024a; Su et al., 2024; Jiao et al., 2024; Su et al., 2025; Zhou et al., 2025; Zhang et al., 2025a; Yang et al., 2025b). These models often exhibit undesirable behaviors, including toxicity, bias, or factual inaccuracies, rooted in the complex and opaque representations learned during training (Yang et al., 2024c; Wang et al., 2025a; Zhang et al., 2025b). Effectively controlling these behaviors without compromising model performance remains an open research problem (Jiao et al., 2025).
(a) Comparison of prior work and MSRS. The left output contains biases and falsehoods. MSRS reduces attribute conflicts by separating steering spaces and using a shared subspace to capture common properties, enabling better integration.
(b) By leveraging both shared and attribute-specific subspaces, MSRS enables effective steering toward attributes $i$ and $j$.
Figure 1: Visualization of MSRS design and comparison with prior work.
@@
Recently, activation steering methods offer a promising avenue for behavior adjustment by manipulating model activations post-training (Im & Li, 2025). Compared to fine-tuning, they offer lightweight control without the need for retraining or access to model weights, enabling scalable adaptation to diverse downstream tasks. These approaches derive an activation steering vector from the difference between the activations of positive and negative samples, applying it during inference to guide outputs toward desired properties without altering model parameters (Rimsky et al., 2024; Zou et al., 2023; Li et al., 2023a; Hu et al., 2024; Yang et al., 2024d; Hong et al., 2024; Yao et al., 2025b; Yu et al., 2025; Yang et al., 2025a). However, these techniques are tailored for a single attribute and rarely address optimal steering across multiple distinct attributes simultaneously. Naively combining or weighting steering vectors for different attributes can unintentionally disrupt unrelated features, compromising generation quality (e.g., fluency or coherence) or inducing conflicts between attribute-specific steering (van der Weij et al., 2024; Ma et al., 2025). For example, enhancing truthfulness may undermine fairness (Wolf et al., 2025) (see Figure 1(a) for an illustration), underscoring a central challenge: mitigating trade-offs to achieve concurrent optimal performance across multiple attributes.
@@
Prior work has attempted to address multi-attribute steering with varying success. For example, ACT (Wang et al., 2025b) employs clustering to train multiple steering probes on positive and negative samples, aiming to capture distinct steering patterns. Similarly, MAT-STEER (Nguyen et al., 2025a) applies orthogonal constraints to activation steering vectors. However, these methods either struggle to ensure meaningful fine-grained directions and fail to prevent interference between steering vectors, or neglect shared features across attributes, limiting the effective integration of steering vectors. The Representation Fine-Tuning (ReFT) (Wu et al., 2024a) based method achieves the goal of steering by fine-tuning model representations in an orthogonal subspace. Orthogonality enables more effective isolation of different attributes at the hidden state level, offering a more principled solution for multi-attribute steering (Zhou, 2025). However, current ReFT approaches face difficulties in subspace allocation, as different attributes demand varying subspace sizes and expressive capacities, which makes their performance suboptimal; simple attributes may require smaller subspaces, while complex ones necessitate larger ones.
@@
To address the poor composability of multiple steering directions, we introduce Multi-Subspace Representation Steering (MSRS), a novel framework that enhances multi-attribute steering through subspace representation fine-tuning, as illustrated in Figure 1(b). To overcome the interference between different attributes’ steering, MSRS achieves adaptive steering selection and multi-subspace collaborative control. Specifically, to reduce interference among attribute-specific directions, MSRS allocates orthogonal subspaces to each attribute, isolating their effects within the representation space. To further tailor subspace capacity to each attribute’s expressive needs, we perform SVD on the attribute-specific activation differences and use leading singular vectors to guide adaptive steering subspace allocation. Finally, MSRS combines attribute-specific subspaces for unique steering directions with an attribute-shared subspace for common steering directions, and learns a dynamic weighting function to compose attribute-specific and shared subspaces efficiently. Furthermore, during inference time, MSRS introduces a dynamic token selection mechanism that identifies and steers the most semantically relevant tokens, enabling token-level intervention and outperforming traditional fixed-position steering approaches.
@@
MSRS demonstrates effectiveness across diverse models (e.g., Llama2-7B (Touvron et al., 2023), Llama3-8B-Instruct (Grattafiori et al., 2024), Qwen2-7B-Instruct (Team, 2024) , Mistral-7B-v0.3 (Jiang et al., 2023)) and tasks (question answering and open-ended generation), significantly reducing attribute conflicts and achieving superior performance across multiple attributes datasets (e.g., concurrent improvements on TruthfulQA(+13%), BBQ(+4%)). Additionally, MSRS generalizes well to standard NLP tasks, achieving gains on HellaSwag (+3.8%) and GLUE (+4.9%). Our contributions can be summarized as follows:
@@
  • During inference, MSRS introduces a dynamic token selection strategy based on subspace similarity, which outperforms the previous fixed token steering.
@@
2 Related Works
Activation Steering Methods. Activation steering aims to adjust activations in specific layers or neurons to guide the model’s output towards desired attributes, without modifying its parameters (Im & Li, 2025). Various approaches have been developed recently (Cao et al., 2024; Bayat et al., 2025; Oozeer et al., 2025). For example, Contrastive Activation Addition (CAA) (Rimsky et al., 2024) computes steering vectors by averaging activation differences between positive and negative examples, which are then added to token positions during inference to control model behavior. Inference-Time Intervention (ITI) (Li et al., 2023a) shifts model activations during inference along predefined directions across attention heads, improving the truthfulness of LLMs. ACT (Wang et al., 2025b) trains multiple steering probes on different steering vectors determined by clustering, obtaining steering vectors for different steering patterns. MAT-Steer uses orthogonal constraints to train activation steering vectors, thereby reducing conflicts between steering directions for different attributes (Nguyen et al., 2025a). However, previous methods primarily address steering for individual attributes or rely on simple combinations of steering vectors. To solve these limitations, we focus on mitigating interference and optimizing composability across multiple attributes.
Representation Fine-Tuning Methods. In these methods, models will be steered through representation editing. Unlike methods that apply one-rank steering vectors, these approaches extend it by using higher-rank matrices and enhance the expressive power of steering vectors, allowing for richer control over model behavior (Wu et al., 2024a). Localized Fine-Tuning (LoFIT) (Yin et al., 2024) identifies critical attention heads for a task and trains offset vectors to modify their hidden representations, offering targeted adjustments. Compositional Subspace Representation Fine-Tuning (CS-ReFT) (Zhou, 2025) advances this by learning orthonormal subspace transformations for distinct skills, composed via a lightweight router, isolating edits in the hidden state to minimize cross-task interference. Unlike previous methods that train steering functions in the same space, we aim to develop representation fine-tuning methods to tune different attribute-specific subspaces and achieve the adaptive integration of multiple attribute steering spaces.
@@ 3 Motivation
To better illustrate our motivation and approach, we first revise ReFT (Wu et al., 2024a). In ReFT, it aims to steer the hidden representation $h\in\mathbb{R}^{d}$ by fine-tuning an $r$-dimensional subspace spanned by the rows of $R$. Specifically, we can define the intervention function $\Phi$ as:
@@
where the learned low-rank projection matrix $R\in\mathbb{R}^{r\times d}$ is typically constrained to have orthonormal rows ($RR^{\top}=I_{r}$), and $W\in\mathbb{R}^{r\times d}$, $b\in\mathbb{R}^{r}$ are trainable parameters. $\Phi(h;R,W,b)$ is integrated into the model’s representations to guide the output towards desired attributes.
The steering-affected output is subsequently optimized to minimize the target objective, thereby refining the parameters $W$, $b$, and $R$. This method enables efficient steering of model representations by manipulating the hidden activations in a learned, low-rank subspace.
However, ReFT demonstrates encounters significant limitations when applied to multi-attribute scenarios. ReFT assumes a single attribute per input, but real-world inputs often involve multiple attributes. Training the matrix $R$ on such multi-attribute inputs forces all steering directions into the same space, causing interference that hinders the model’s ability to balance the needs of each attribute, ultimately limiting its performance across all attributes.  Zhou (2025) attempts to address this by partitioning $R$ into equal-sized subspaces, each dedicated to a specific attribute, in an effort to reduce interference. However, this approach overlooks the fact that different attributes require subspaces of varying sizes based on their expressive needs. As a result, attributes with higher complexity may not receive enough capacity for effective steering, while simpler attributes may waste valuable space.
@@ 4.1 Multi-Attribute Steering Direction Extraction
Attribute-wise Activation Aggregation. To extract the primary steering directions for each attribute from the activation values, we first capture the key feature representations for each attribute. Specifically, for each attribute $i$, we compute the average activation $\tau_{i}$ from its corresponding dataset $\mathcal{D}_{i}$. We extract the model’s intermediate activation $h_{i,j}^{l}$ of layer $l$ for each sample $j$ at the last few tokens, specifically averaging the activations from these tokens, as they capture the full context of the prompt  (Lei & Cooper, 2025). The average activation for attribute $i$ is then computed as $\tau_{i}=\frac{1}{|\mathcal{D}_{i}|}\sum_{j=1}^{|\mathcal{D}_{i}|}h_{i,j}^{l}$. To integrate information across all $n$ attributes, we construct a combined activation matrix $\tau_{c}=[\tau_{1}\;|\;\tau_{2}\;|\;\dots\;|\;\tau_{n}]\in\mathbb{R}^{d\times n}$.
@@
Shared and Specific Subspace Extraction. To retain common knowledge while enabling attribute-specific steering, we perform singular value decomposition (SVD) on the aggregated activation matrix $\tau_{c}$, i.e., $\tau_{c}=U_{c}\Sigma_{c}V_{c}^{\top}$. We adaptively select the smallest number $r_{s}$ such that the cumulative energy (sum of top $r_{s}$ singular values) accounts for at least 60% of the total energy in $\Sigma_{c}$. This yields the shared subspace of $V_{c}$, defined as $B_{\text{shared}}=V_{c,1:r_{s}}^{\top}\in\mathbb{R}^{r_{s}\times d}$. Intuitively, $B_{\text{shared}}$ captures the dominant shared directions across all attributes. By selecting the top singular vectors, we capture high-variance directions that represent the most expressive steering dimensions. It allows us to automatically allocate varying subspace sizes for each attribute based on its expressive needs. For complex attributes, it selects more top vectors, while simpler attributes are allocated smaller subspaces, effectively addressing the mismatch between each attribute’s steering capacity and its allocated subspace size.
@@
For each attribute $i$, we then isolate attribute-specific directions by projecting $\tau_{i}$ onto the shared subspace and computing the residual, i.e., $H_{\text{res}}^{(i)}=\tau_{i}-B_{\text{shared}}^{\top}B_{\text{shared}}\tau_{i}$. Applying SVD to $H_{\text{res}}^{(i)}$, we obtain $H_{\text{res}}^{(i)}=U^{(i)}S^{(i)}V^{(i)\top}$. Similarly, we select the smallest number $r_{i}$ such that the top $r_{i}$ singular values of $S^{(i)}$, and define the private subspace as $B_{i}=\left(V^{(i)}_{1:r_{i}}\right)^{\top}\in\mathbb{R}^{r_{i}\times d}$. Generally, $B_{i}$ captures directions orthogonal to the shared subspace, preserving attribute-specific semantics. This adaptive allocation allows each attribute subspace to retain only as much representational capacity as needed, reflecting its inherent complexity.
@@ 4.2 Adaptive Subspace Selecting
Based on equation 1, we introduce a mask network $m(h)=\text{sigmoid}(\text{MLP}(h))\in[0,1]^{r}$, which assigns weights to each subspace dimension. The intervention function becomes:

### v4

# Adaptive Multi-Subspace Representation Steering for Attribute Alignment in Large Language Models

@@ Abstract
Activation steering offers a promising approach to controlling the behavior of Large Language Models by directly manipulating their internal activations. However, most existing methods struggle to jointly steer multiple attributes, often resulting in interference and undesirable trade-offs. To address this challenge, we propose Multi-Subspace Representation Steering (MSRS), a novel framework for effective multi-attribute steering via subspace representation fine-tuning. MSRS reduces inter-attribute interference by allocating orthogonal subspaces to each attribute, isolating their influence within the model’s representation space. MSRS also incorporates a hybrid subspace composition strategy: it combines attribute-specific subspaces for unique steering directions with a shared subspace for common steering directions. A dynamic weighting function learns to efficiently integrate these components for precise control. Experimental results show that MSRS significantly reduces attribute conflicts, surpasses existing methods across a range of attributes, and generalizes effectively to diverse downstream tasks.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing, driving advancements in applications such as text generation, question answering, and dialogue systems (Qin et al., 2024; Matarazzo and Torlone, 2025). However, as LLMs are increasingly deployed in real-world, sensitive contexts, ensuring their behavior aligns with desired attributes, such as truthfulness and fairness, has become a critical challenge (Yang et al., 2024; Su et al., 2024; Jiao et al., 2024). These models often exhibit undesirable behaviors, including toxicity, bias, or factual inaccuracies, rooted in the complex and opaque representations learned during training (Le Bronnec et al., 2024). Effectively controlling these behaviors without compromising model performance remains an open research problem (Jiao et al., 2025).
Figure 1: Comparison of prior work and MSRS.
@@
Recently, activation steering methods offer a promising avenue for behavior adjustment by manipulating model activations post-training (Im and Li, 2025). Compared to fine-tuning, they offer lightweight control without the need for retraining or access to model weights, enabling scalable adaptation to diverse downstream tasks. These approaches derive an activation steering vector from the difference between the activations of positive and negative samples, applying it during inference to guide outputs toward desired properties without altering model parameters (Rimsky et al., 2024; Zou et al., 2023; Li et al., 2023a). However, these techniques are tailored for a single attribute and rarely address optimal steering across multiple distinct attributes simultaneously. Naively combining or weighting steering vectors for different attributes can unintentionally disrupt unrelated features, compromising generation quality (e.g., fluency or coherence) or inducing conflicts between attribute-specific steering (van der Weij et al., 2024; Ma et al., 2025). For example, enhancing truthfulness may undermine fairness (Wolf et al., ) (see Figure 1 for an illustration), underscoring a central challenge: mitigating trade-offs to achieve concurrent optimal performance across multiple attributes.
@@
Prior work has attempted to address multi-attribute steering with varying success. For example, ACT (Wang et al., 2025) employs clustering to train multiple steering probes on positive and negative samples, aiming to capture distinct steering patterns. Similarly, MAT-STEER (Nguyen et al., 2025) applies orthogonal constraints to activation steering vectors. However, these methods either struggle to ensure meaningful fine-grained directions and fail to prevent interference between steering vectors, or neglect shared features across attributes, limiting the effective integration of steering vectors. The Representation Fine-Tuning (ReFT) (Wu et al., 2024) based method achieves the goal of steering by fine-tuning model representations in an orthogonal subspace. Orthogonality enables more effective isolation of different attributes at the hidden state level, offering a more principled solution for multi-attribute steering (Zhou, 2025). However, it faces difficulties in subspace allocation, as different attributes demand varying subspace sizes and expressive capacities, which makes their performance suboptimal; simple attributes may require smaller subspaces, while complex ones necessitate larger ones.
@@
To address the poor composability of multiple steering directions, we introduce Multi-Subspace Representation Steering (MSRS), a novel framework that enhances multi-attribute steering through subspace representation fine-tuning, as illustrated in Figure 2. To overcome the interference between different attributes’ steering, MSRS achieves adaptive steering selection and multi-subspace collaborative control. Specifically, to reduce interference among attribute-specific directions, MSRS allocates orthogonal subspaces to each attribute, isolating their effects within the representation space. To further tailor subspace capacity to each attribute’s expressive needs, we perform SVD on the attribute-specific activation differences and use leading singular vectors to guide adaptive steering subspace allocation. Finally, MSRS combines attribute-specific subspaces for unique steering directions with an attribute-shared subspace for common steering directions, and learns a dynamic weighting function to compose attribute-specific and shared subspaces efficiently.
@@
MSRS demonstrates effectiveness across diverse models (e.g., Llama2-7B (Touvron et al., 2023), Llama3-8B-Instruct (Grattafiori et al., 2024), Qwen2-7B-Instruct (Team, 2024) , Mistral-7B-v0.3 (Jiang et al., 2023)) and tasks (multiple-choice and open-ended generation), significantly reducing attribute conflicts and achieving superior performance across multiple attributes datasets (e.g., concurrent improvements on TruthfulQA(+13%), BBQ(+4%)). Additionally, MSRS generalizes well to standard NLP tasks, achieving gains on HellaSwag (+3.8%) and GLUE (+4.9%). Our contributions can be summarized as follows:
@@
  • MSRS demonstrates effective scalability across multiple attributes, maintaining performance gains in practical multi-attribute scenarios.
@@
2 Related Work
Activation Steering Methods. Activation steering guides model behavior by adjusting internal activations without modifying parameters (Im and Li, 2025). Early approaches like CAA (Rimsky et al., 2024) and ITI (Li et al., 2023a) rely on static vectors derived from contrastive pairs or specific attention heads. While recent works like ACT (Wang et al., 2025) and MAT-Steer (Nguyen et al., 2025) incorporate clustering or orthogonal constraints to handle multiple patterns, they primarily focus on individual attributes or simple vector combinations. However, previous methods primarily address steering for individual attributes or rely on simple combinations of steering vectors. To solve these limitations, we focus on mitigating interference and optimizing composability across multiple attributes.
Representation Fine-Tuning Methods. Unlike single-vector steering, representation fine-tuning learns higher-rank transformations to enhance control expressivity (Wu et al., 2024). Methods such as LoFIT (Yin et al., 2024) offer targeted adjustments by modifying critical attention heads. More recently, CS-ReFT (Zhou, 2025) advanced this by learning orthonormal subspaces for distinct skills composed via a router. However, while CS-ReFT emphasizes task isolation via routing, our approach focuses on the adaptive integration of attribute-specific subspaces. Unlike previous methods that train steering functions in the same space, we aim to develop representation fine-tuning methods to tune different attribute-specific subspaces and achieve the adaptive integration of multiple attribute steering spaces. See more details in Appendix J.
@@ 3 Motivation
To better illustrate our motivation and approach, we first revise ReFT (Wu et al., 2024). In ReFT, it aims to steer the hidden representation $h\in\mathbb{R}^{d}$ by fine-tuning an $r$-dimensional subspace spanned by the rows of $R$. Specifically, we can define the intervention function $\Phi$ as:
@@
where the learned low-rank projection matrix $R\in\mathbb{R}^{r\times d}$, is typically constrained to have orthonormal rows ($RR^{\top}=I_{r}$), and $W\in\mathbb{R}^{r\times d}$, $b\in\mathbb{R}^{r}$ are trainable parameters. $\Phi(h;R,W,b)$ is integrated into the model’s representations to guide the output towards desired attributes. The steering-affected output is subsequently optimized to minimize the target objective, thereby refining the parameters $W$, $b$, and $R$. This method enables efficient steering of model representations by manipulating the hidden activations in a learned, low-rank subspace.
While ReFT encounters significant limitations when applied to multi-attribute scenarios. ReFT assumes a single attribute per input, but real-world inputs often involve multiple attributes. Training the matrix $R$ on such multi-attribute inputs forces all steering directions into the same space, causing interference that hinders the model’s ability to balance the needs of each attribute, ultimately limiting its performance across all attributes.  Zhou (2025) attempts to address this by partitioning $R$ into equal-sized subspaces, each dedicated to a specific attribute, in an effort to reduce interference. However, this approach overlooks the fact that different attributes require subspaces of varying sizes based on their expressive needs. As a result, attributes with higher complexity may not receive enough capacity for effective steering, while simpler attributes may waste valuable space.
@@ 4 Methodology
Figure 2: By leveraging both shared and attribute-specific subspaces, MSRS enables effective steering toward attributes $i$ and $j$.
@@ 4.1 Multi-Attribute Steering Direction Extraction
Attribute-wise Activation Aggregation. To extract the primary steering directions for each attribute from the activation values, we first capture the key feature representations for each attribute. Specifically, for each attribute $i$, we compute the average activation $\tau_{i}$ from its corresponding dataset $\mathcal{D}_{i}$. We extract the model’s intermediate activation $h_{i,j}^{l}$ of layer $l$ for each sample $j$ at the last few tokens, specifically averaging the activations from these tokens, as they capture the full context of the prompt  (Lei and Cooper, 2025). The average activation for attribute $i$ is then computed as $\tau_{i}=\frac{1}{|\mathcal{D}_{i}|}\sum_{j=1}^{|\mathcal{D}_{i}|}h_{i,j}^{l}$. To integrate information across all $n$ attributes, we construct a combined activation matrix $\tau_{c}=[\tau_{1}\;|\;\tau_{2}\;|\;\dots\;|\;\tau_{n}]\in\mathbb{R}^{d\times n}$.
@@
Shared and Specific Subspace Extraction. To retain common knowledge while enabling attribute-specific steering, we perform singular value decomposition (SVD) on the aggregated activation matrix $\tau_{c}$, i.e., $\tau_{c}=U_{c}\Sigma_{c}V_{c}^{\top}$. We adaptively select the smallest number $r_{s}$ such that the cumulative energy (sum of top $r_{s}$ singular values) accounts for at least 60% of the total energy in $\Sigma_{c}$. This yields the shared subspace of $U_{c}$, defined as $B_{\text{shared}}=U_{c,1:r_{s}}^{\top}\in\mathbb{R}^{r_{s}\times d}$. Intuitively, $B_{\text{shared}}$ captures the dominant shared directions across all attributes.
For each attribute $i$, we then iteratively isolate attribute-specific directions by projecting $H_{i}^{(i)}$ onto both the shared subspace and all previously extracted private subspaces, computing the residual, i.e.,
$H_{\text{res}}^{(i)}=H_{i}^{(i)}-B_{\text{shared}}^{\top}B_{\text{shared}}H_{i}^{(i)}-\sum_{k=1}^{i-1}B_{k}^{\top}B_{k}H_{i}^{(i)},$ (2)
@@
where $H_{i}^{(i)}$ denotes the activation matrix for attribute $i$, formed by concatenating the sample representations, i.e., $H_{i}^{(i)}=\left[h_{i,1}^{l},h_{i,2}^{l},\dots,h_{i,D_{i}}^{l}\right]$. Applying SVD to $H_{\text{res}}^{(i)}$, we obtain $H_{\text{res}}^{(i)}=U^{(i)}S^{(i)}V^{(i)\top}$. Similarly, we select the smallest number $r_{i}$ such that the top $r_{i}$ singular values of $S^{(i)}$ capture the principal energy, and define the private subspace as $B_{i}=\left(U^{(i)}_{1:r_{i}}\right)^{\top}\in\mathbb{R}^{r_{i}\times d}$. Generally, $B_{i}$ captures directions orthogonal to the shared subspace and prior attributes, preserving strictly unique attribute-specific semantics. By selecting the top singular vectors, we capture high-variance directions that represent the most expressive steering dimensions. It allows us to automatically allocate varying subspace sizes for each attribute based on its expressive needs. For complex attributes, it selects more top vectors, while simpler attributes are allocated smaller subspaces, effectively addressing the mismatch between each attribute’s steering capacity and its allocated subspace size. This adaptive allocation allows each attribute subspace to retain only as much representational capacity as needed, reflecting its inherent complexity.
@@ 4.2 Adaptive Subspace Selecting
Based on (1), we introduce a mask network $m(h)=\text{sigmoid}(\text{MLP}(h))\in[0,1]^{r}$, which assigns weights to each subspace dimension. The intervention function becomes: